OpenStudy (anonymous):

i need help converting powers to decimals 41 to the 2nd power

jimthompson5910 (jim_thompson5910):

\[\large 41^2=41*41 = 1681\] So \[\large 41^2 = 1681\]

OpenStudy (anonymous):

so you dont put a decimal in the answer even though they say to?

jimthompson5910 (jim_thompson5910):

you can write 1681.0, but that's not necessary

OpenStudy (anonymous):

is there a site on line with calculators to do this kind of math?

jimthompson5910 (jim_thompson5910):

google is a calculator if you type in math expressions wolfram alpha is also a good calculator: http://www.wolframalpha.com

OpenStudy (anonymous):

Thanks for the help.

OpenStudy (anonymous):

how do you make the small number the exponent number

jimthompson5910 (jim_thompson5910):

To type something like \[2^3\] type 2^3